# Puzzle: Distillation-Based NAS for Inference-Optimized LLMs

### v3

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities, but their adoption is limited by high computational costs during inference. While increasing parameter counts enhances accuracy, it also widens the gap between state-of-the-art capabilities and practical deployability. We present Puzzle, a framework to accelerate LLM inference on specific hardware while preserving their capabilities. Through an innovative application of neural architecture search (NAS) at an unprecedented scale, Puzzle systematically optimizes models with tens of billions of parameters under hardware constraints. Our approach utilizes blockwise local knowledge distillation (BLD) for parallel architecture exploration and employs mixed-integer programming for precise constraint optimization.
We demonstrate the real-world impact of our framework through Llama-3.1-Nemotron-51B-Instruct (Nemotron-51B), a publicly available model derived from Llama-3.1-70B-Instruct. Nemotron-51B achieves a 2.17x inference throughput speedup, fitting on a single NVIDIA H100 GPU while preserving 98.4% of the original model’s capabilities. Nemotron-51B currently stands as the most accurate language model capable of inference on a single GPU with large batch sizes. Remarkably, this transformation required just 45B training tokens, compared to over 15T tokens used for the 70B model it was derived from. This establishes a new paradigm where powerful models can be optimized for efficient deployment with only negligible compromise of their capabilities, demonstrating that inference performance, not parameter count alone, should guide model selection. With the release of Nemotron-51B and the presentation of the Puzzle framework, we provide practitioners immediate access to state-of-the-art language modeling capabilities at significantly reduced computational costs.
@@ 2 Search Space
The motivation for applying NAS in our work lies in the redundancy found in many trained LLMs. Numerous studies have shown that a significant portion of computations and parameters in LLMs become redundant post-training, during inference [44, 24, 3]. Prior studies tried to solve these issues with techniques such as pruning [38, 35, 7, 34, 49, 48], removing entire layers [18], local attention methods (e.g., window and sink attention) [9, 50], reducing the number of key-value (KV) heads [45, 4] and many more. Given the high computational and monetary cost associated with running LLMs, optimizing these models to eliminate inefficiencies becomes a critical goal. NAS methods are defined by the search space, the search strategy and evaluation metric of candidate architectures. NAS offers a systematic approach to exploring architectural changes that balances performance and resource constraints, and thus is a prime candidate for reducing inference costs.
@@ 3.1 Building a Block Library with Decoupled Blockwise Local Distillation
The first stage of our decomposed NAS framework (further discussed in Section 4) is building a “library” of trained blocks. In order to cover the entire search space defined in Section 2, we need to obtain trained weights for each attention variant $a_{j}$ and each FFN variant $f_{k}$ in each transformer layer $i$. We consider two methods to train the block library: coupled BLD and decoupled BLD. For each transformer layer $i$, coupled BLD constructs each possible block variant $[a_{j},f_{k}]_{i}$ and trains it to emulate the corresponding parent block $[a_{\text{parent}},f_{\text{parent}}]_{i}$. This approach is similar to the aforementioned decomposed NAS methods used in CV. Given the significantly higher computational costs of LLMs compared to CV models, and noting the inherent structure of the transformer layer, we propose decoupled BLD to drastically reduce the cost of building a block library: training $[a_{j},f_{\text{parent}}^{(\text{frozen})}]_{i}$ and $[a_{\text{parent}}^{(\text{frozen})},f_{k}]_{i}$ separately to emulate $[a_{\text{parent}},f_{\text{parent}}]_{i}$ while freezing the weights of the student subblock that is identical to the parent, and composing the trained subblocks into a full block $[a_{j},f_{k}]_{i}$ only after training.
@@ 4.1 Estimating Resource Requirements
Memory requirements during inference comprise two distinct components with different scaling behaviors. The parameter memory, while substantial, remains constant regardless of the input size. In contrast, the key-value cache memory scales linearly with both batch size and sequence length, often becoming the dominant factor in long-sequence scenarios. For example, in a model with 32 attention heads and 128-dimensional head size, each token requires 8KB of KV-cache per layer using FP16 precision. For an 8K-token sequence with batch size 64, this amounts to 4GB per layer just for the KV-cache, potentially exceeding the parameter memory.
@@ 4.2 Scoring Architecture Solutions
We score each block variant at each network location by measuring the impact of replacing only that specific block in the parent model. To do this, we construct a model identical to the parent but with a single block replaced by a trained block from our library, then calculate a performance measure on the entire model.For efficient I/O, when scoring multiple variants, we load onto the GPU only the blocks that differ from the previously evaluated model. We call these scores replace-1-block scores. During architecture search, we estimate the quality of a constructed architecture as the sum of its individual replace-1-block scores (see Section 4.3). Note that the scoring is not performed on candidate models during the search phase – we estimate the quality of each block only once, then use its replace-1-block score to estimate the quality of any candidate that contains it.
@@ 7 Main Results
Human evaluation: We complement the above accuracy performance benchmarks with a human evaluation comparing the parent and child models. A blind test comparison of the models was conducted on a general purpose test set that includes tasks such as reasoning, longform text generation, knowledge and more. Results show that the difference between the models is not statistically significant (Figure 4), strengthening the claim that the accuracy degradation is minimal. For more details about the evaluation see appendix A.1.
@@
Figure 5: Accuracy vs. Throughput performance of Nemotron-51B compared to state-of-the-art models. Throughput is measured on NVIDIA H100 GPUs with the optimal TP setting per model, all running in FP8 on a “text generation” scenario (see Table 3). The red line represents the efficient frontier, highlighting models with the best accuracy-to-throughput tradeoff. Accuracy=(MT-Bench $\times$10 + MMLU) / 2
@@
Evaluating Long-Context Performance: To evaluate the long-context performance of Nemotron-51B, we employed a subset of the RULER benchmark suite  [23]. Llama-3.1-70B-Instruct, with a maximum context length of 128K tokens, serves as the performance upper bound. Notably, the child model was only exposed to sequences up to 8K tokens during training. Despite this limitation, the child model exhibited remarkable and somewhat surprising performance, retaining 96.23% of its parent’s accuracy at 16K tokens across the RULER benchmark. This outcome highlights the robustness of Puzzle’s framework in enabling generalization beyond the direct training horizon. Specifically, it is noteworthy that a child model trained solely on sequences up to 8K tokens could maintain such high accuracy at double the context length. We hypothesize that a short fine-tuning phase on longer contexts could further extend the model’s effective context length.
At 32K tokens, the child model maintained a significant portion of the teacher’s capabilities before performance degraded at 64K and beyond. This result aligns with expectations, given the child model’s exposure constraints during training. These findings highlight the potential for efficient training strategies to extend the effective context length of language models beyond their direct training horizon.
Table 4 compares the performance of the parent and child models on benchmarks up to 16K tokens. Full performance details for extended context lengths are provided in Appendix B.
Table 4: Performance comparison of Llama-3.1-70B-Instruct (parent) and Nemotron-51B (child) on the RULER benchmark for context lengths up to 16K tokens. Accuracy preserved is the ratio of the Nemotron average to the Llama average, expressed as a percentage.
Context Length Parent Average Score Child Average Score Accuracy Preserved (%)
1024 99.81 99.78 99.90
2048 97.97 97.32 99.34
4096 93.38 92.12 98.65
8192 92.68 90.63 97.79
16384 91.08 87.65 96.23
Finally, we applied the Puzzle framework to produce a child derivative of Llama-3.1-8B-Instruct, optimized for throughput specifically on an NVIDIA RTX 4090 GPU. This model breaks the efficient frontier for its throughput range, demonstrating Puzzle’s ability to deliver highly efficient architectures also on consumer-grade hardware, while preserving the balance between accuracy and performance. Table 5 highlights this model’s superior tradeoff in accuracy and efficiency.
@@ 8.1.1 Combining Coupled BLD and Decoupled BLD
Puzzle with Decouple BLD 5834 72.92
@@ A.1 Human Evaluation
Appendix B RULER Benchmark Performance Tables
This section provides the complete performance tables for both the parent (Llama-3.1-70B-Instruct) and child (Nemotron-51B) models on a subset of the RULER benchmark across all context lengths evaluated.

### v4

@@ Abstract
Large language models (LLMs) offer remarkable capabilities, yet their high inference costs restrict wider adoption. While increasing parameter counts improves accuracy, it also broadens the gap between state-of-the-art capabilities and practical deployability. We present Puzzle, a hardware-aware framework that accelerates the inference of LLMs while preserving their capabilities. Using neural architecture search (NAS) at a large-scale, Puzzle optimizes models with tens of billions of parameters. Our approach utilizes blockwise local knowledge distillation (BLD) for parallel architecture exploration and employs mixed-integer programming for precise constraint optimization.
We showcase our framework’s impact via Llama-3.1-Nemotron-51B-Instruct (Nemotron-51B), a publicly available model derived from Llama-3.1-70B-Instruct. Nemotron-51B achieves a 2.17$\times$ inference throughput speedup, fitting on a single NVIDIA H100 GPU while retaining 98.4% of the original model’s benchmark accuracies. Notably, it is the most accurate model supporting single H100 GPU inference with large batch sizes, despite training on only 45B tokens, far fewer than the 15T used to train Llama-70B. Lastly, we derive Llama-3.3-Nemotron-49B-Super-Base to demonstrate Puzzle can retain long-context and that lightweight alignment on these derived models allows them to surpass the parent model in specific capabilities. Our work establishes that powerful LLM models can be optimized for efficient deployment with only negligible loss in quality, underscoring that inference performance, not parameter count alone, should guide model selection.
@@ 2 Search Space
The motivation for applying NAS in our work lies in the redundancy found in many trained LLMs. Numerous studies have shown that a significant portion of computations and parameters in LLMs become redundant post-training, during inference [46, 25, 3]. Prior studies tried to solve these issues with techniques such as pruning [40, 37, 7, 36, 52, 51], removing entire layers [19, 14], local attention methods (e.g., window and sink attention) [9, 53], reducing the number of key-value (KV) heads [47, 4] and many more. Given the high computational and monetary cost associated with running LLMs, optimizing these models to eliminate inefficiencies becomes a critical goal. NAS methods are defined by the search space, the search strategy and evaluation metric of candidate architectures. NAS offers a systematic approach to exploring architectural changes that balances performance and resource constraints, and thus is a prime candidate for reducing inference costs.
@@ 3.1 Building a Block Library with Decoupled Blockwise Local Distillation
The first stage of our decomposed NAS framework (further discussed in Section 4) is building a “library” of trained blocks. In order to cover the entire search space defined in Section 2, we need to obtain trained weights for each attention variant $a_{j}$ and each FFN variant $f_{k}$ in each transformer layer $i$. We consider two methods to train the block library: coupled BLD and decoupled BLD. For each transformer layer $i$, coupled BLD constructs each possible block variant $[a_{j},f_{k}]_{i}$ and trains it to emulate the corresponding parent block $[a_{\text{parent}},f_{\text{parent}}]_{i}$. This approach is similar to the aforementioned decomposed NAS methods used in CV. Given the significantly higher computational costs of LLMs compared to CV models, and noting the inherent structure of the transformer layer, we propose decoupled BLD to drastically reduce the cost of building a block library: training $[a_{j},f_{\text{parent}}^{(\text{frozen})}]_{i}$ and $[a_{\text{parent}}^{(\text{frozen})},f_{k}]_{i}$ separately to emulate $[a_{\text{parent}},f_{\text{parent}}]_{i}$ while freezing the weights of the child subblock that is identical to the parent, and composing the trained subblocks into a full block $[a_{j},f_{k}]_{i}$ only after training.
@@ 4.1 Estimating Resource Requirements
Memory requirements during inference comprise two distinct components with different scaling behaviors. Parameter memory, while substantial, remains constant regardless of the input size. In contrast, the key-value cache memory scales linearly with both batch size and sequence length, often becoming the dominant factor in long-sequence scenarios. For example, in a model with 32 attention heads and 128-dimensional head size, each token requires 8KB of KV-cache per layer using FP16 precision. For an 8K-token sequence with batch size 64, this amounts to 4GB per layer just for the KV-cache, potentially exceeding the parameter memory.
@@ 4.2 Scoring Architecture Solutions
We score each block variant at each network location by measuring the impact of replacing only that specific block in the parent model. To do this, we construct a model identical to the parent but with a single block replaced by a trained block from our library, then calculate a performance measure on the entire model. For efficient I/O, when scoring multiple variants, we load onto the GPU only the blocks that differ from the previously evaluated model. We call these scores replace-1-block scores. During architecture search, we estimate the quality of a constructed architecture as the sum of its individual replace-1-block scores (see Section 4.3). Note that scoring is not performed on candidate models during the search phase – we estimate the quality of each block only once, then use its replace-1-block score to estimate the quality of any candidate that contains it.
@@ 7 Main Results
Human evaluation: We complement the above accuracy performance benchmarks with a human evaluation comparing the parent and child models. A blind test comparison of the models was conducted on a general purpose test set that includes tasks such as reasoning, longform text generation, knowledge and more. Results show comparable performance between the models (Figure 4), strengthening the claim that the accuracy degradation is minimal. For more details about the evaluation see Appendix A.1.
@@
Figure 5: Accuracy vs. Throughput performance of Nemotron-51B compared to state-of-the-art models. Throughput is measured on NVIDIA H100 GPUs with optimal TP setting per model, all running in FP8 on a “text generation” scenario (see Table 3). The red line represents the efficient frontier, highlighting models with the best accuracy-to-throughput tradeoff. Accuracy=(MT-Bench $\times$10 + MMLU) / 2
@@
Long-Context Performance: We evaluated Nemotron-51B on a subset of the RULER benchmark [24]. Notably, although Nemotron-51B was trained only on sequences up to 8K tokens, it retained over 96% of its parent’s accuracy at 16K tokens, highlighting Puzzle’s ability to preserve performance beyond the direct training length. Performance degraded beyond 64K tokens, suggesting that fine tuning on longer contexts could extend the effective context length.
To demonstrate this extension, we produced Llama-3.3-Nemotron-49B-Super-Base (henceforth Nemotron-49B-Base, the base version of Llama-3.3-Nemotron-49B-Superhttps://huggingface.co/nvidia/Llama-3_3-Nemotron-Super-49B-v1) from Llama-3.3-70B-Instruct with identical constraints to Nemotron-51B and uptrained on an additional 5B tokens at 64K context and 5B at 128K. Table 4 shows that Nemotron-49B-Base maintains or exceeds its parent’s performance up to 16K tokens, retains over 98% at 64K, and remains above 94% at 128K. Full performance details are provided in Appendix B.
Table 4: Comparison of Llama-3.3-70B-Instruct (parent) and Nemotron-49B-Base (child) on RULER for context lengths up to 128K.
Context Parent Average Score Child Average Score Accuracy Preserved (%)
4K 96.77 97.40 100.65
8K 96.46 96.59 100.13
16K 95.98 96.09 100.11
32K 94.70 94.30 99.57
64K 88.91 87.39 98.29
128K 52.25 49.62 94.96
Alignment: We further aligned Nemotron-49B-Base by following the RLHF and instruction-tuning recipe of Wang et al. [49], using 25M tokens. Table 5 compares results before and after alignment, showing that Puzzle derivatives can seamlessly undergo RLHF-based alignment to boost their accuracy.
Table 5: The performance of Nemotron-49B-Base before and after alignment, even surpassing its parent in certain benchmarks.
Model MMLU MT-Bench Arena Hard [32]
Nemotron-49B-Base (after lightweight alignment) 80.98 9.21 82.1
Nemotron-49B-Base (before alignment) 81.03 8.77 65.8
Llama-3.3-70B-Instruct (parent) 81.66 8.84 71.70
Compact model: we applied the Puzzle framework to produce a child derivative of Llama-3.1-8B-Instruct, optimized for throughput specifically on an NVIDIA RTX 4090 GPU. This model breaks the efficient frontier for its throughput range, demonstrating Puzzle’s ability to deliver highly efficient architectures also on consumer-grade hardware, while preserving the balance between accuracy and performance. Table 6 highlights this model’s superior tradeoff in accuracy and efficiency.
@@ 8.1.1 Combining Coupled BLD and Decoupled BLD
Puzzle with Decouple BLD 5834 73.10
@@ A.1 Human Evaluation
Appendix B RULER Benchmark Performance Tables
This section provides complete performance tables for our parent-child pairs on a subset of the RULER benchmark across all context lengths evaluated.
B.1 Nemotron-51B vs. Llama-3.1-70B-Instruct
Table 15 shows the results of the parent (Llama-3.1-70B-Instruct) and child (Nemotron-51B) for all context lengths. As noted in the main text, Nemotron-51B was trained on sequences up to only 8K tokens yet retains more than 96% of its parent’s performance at 16K. The child’s performance degrades at 64K and beyond, which is unsurprising given its training horizon. Nevertheless, these results underscore that a large fraction of the parent’s long-context capabilities can remain intact even without explicit training on such long sequences.
B.2 Nemotron-49B-Base vs. Llama-3.3-70B-Instruct
In Table 16, we present the extended context-length results for Nemotron-49B-Base (uptrained with sequences up to 128K tokens) alongside its parent (Llama-3.3-70B-Instruct). We rename columns to make them more consistent with the style above, although some tasks differ from those used for Nemotron-51B. Nemotron-49B-Base preserves 98% or more of its parent’s performance up to 64K tokens and remains above 94% at 128K. This highlights that adding a short uptraining phase on longer contexts can effectively extend the context range of Puzzle-derived models. Additional details and per-task results appear below.
Table 16: Performance comparison of the parent (Llama-3.3-70B-Instruct) and child (Nemotron-49B-Base) on a subset of the RULER benchmark. ‘Accuracy Preserved’ is (child / parent)$\times$100.
Context single_needle_1 single_needle_2 single_needle_3 multi_needle_1 multi_needle_2 multi_needle_3 multi_value multi_query variable_tracking_1_4 common_words_extraction freq_words_extraction qa_squad qa_hotpotqa Average Accuracy Preserved (%)
Parent: Llama-3.3-70B-Instruct
4K 100 100 100 100 100 100 100 100 100 95 91 72 96.77 96.77 -
8K 100 100 100 100 100 100 100 100 100 96 87 71 96.46 96.46 -
16K 100 100 100 100 100 100 100 100 99.8 95 86 67 95.98 95.98 -
32K 100 100 100 100 100 99 98.75 100 95 92.33 83 63 94.70 94.70 -
64K 100 100 100 100 97 93 98.25 100 43.9 92.67 75 56 88.91 88.91 -
128K 37 100 100 90 0 1 98.25 95.75 0.4 3.5 78.33 41 34 52.25 52.25 -
Child: Nemotron-49B-Base
4K 100 100 100 100 100 100 100 100 100 95.9 99.33 93 78 97.40 100.65
8K 100 100 100 100 100 100 100 100 100 92.4 98.33 89 76 96.59 100.13
16K 100 100 100 100 99 100 99.75 100 100 92.4 98 89 71 96.09 100.11
32K 100 100 100 100 99 99 99.5 100 81.4 95 85 67 94.3 94.30 99.58
64K 100 100 100 100 92 92 99.75 100 26.9 92.67 73 60 87.39 87.39 98.29
128K 45 99 99 89 0 0 88 86.5 1.6 1.3 56.67 44 35 49.62 94.97